OpenStudy (anonymous):

Which pair of ratios forms a proportion? 10 over 22. and 5 over 44. 16 over 9. and 8 over 18. 7 over 8. and 14 over 16. 20 over 24. and 40 over 12.

ganeshie8 (ganeshie8):

10 over 22. and 5 over 44. \(\large \frac{10}{22} = \frac{5}{44}\) cross multiply and see if they match

ganeshie8 (ganeshie8):

\(\large \frac{10}{22} = \frac{5}{44}\) \(\large 10 \times 44 = 5 \times 22\) \(\large 440 = 110 \) FALSE, so REJECT, test next option :)

OpenStudy (anonymous):

thanks

ganeshie8 (ganeshie8):

np :) have u found the correct proportion ?